An electric light bulb has a volume of 200cm3, if contains argon gas at a temperature of 1.1×10 and a temperature of 25°C. When the light is switched on, the pressure increases to a steady 1.8 ×10 Pa. What is the temperature of the argon in the electric light bulb that creates this new steady pressure?
1
Expert's answer
2020-04-02T08:56:13-0400

V = const = 200 cm3

T1 = 250C + 273 = 298 K

P1 = 1.1x10... Pa


P2 = 1.8 ×10... Pa

T2 = P2*T1/P1 = 1.8*298/1.1 = 487.5 K

t1 = 487.5 - 273 = 214.5 0C